OpenStudy (anonymous):

Use matrices to solve the system of equations. (If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.) SYSTEM IS IN THE COMMENTS>

OpenStudy (anonymous):

2x+3y-2z=16 5x-6y+z=22 4y-2z=4

OpenStudy (haleyelizabeth2017):

Do you know how to set it up?

OpenStudy (haleyelizabeth2017):

 Okay?